On April 2, 2017, Montana Mining Co. pays $4,159,740 for an ore deposit containing 1,516,000 tons. The company installs machinery in the mine costing $216,200, with an estimated seven-year life and no salvage value. The machinery will be abandoned when the ore is completely mined. Montana begins mining on May 1, 2017, and mines and sells 180,400 tons of ore during the remaining eight months of 2017.
  
Prepare the December 31, 2017, entries to record both the ore deposit depletion and the mining machinery depreciation. Mining machinery depreciation should be in proportion to the mine’s depletion. (Do not round intermediate calculations. Round your final answers to the nearest whole number.)
Answer #1

Deposit depletion = (4159740/1516000) *180400 = $494998

Depreciation on machine = (216200/1516000) *180400 =$25727

Accounts Debit Credit
Depletion expense - Ore $494, 998
Accumulated depletion - Ore $494, 998
Depreciation expense $25727
Accumulated depreciation $25727